OpenStudy (anonymous):

use the square root property to solve the equation (5x+5)(exp. 2)=100

OpenStudy (jennychan12):

square root both sides then solve for x 

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

Yes :)

OpenStudy (openstudier):

Hold on!

OpenStudy (anonymous):

k

OpenStudy (openstudier):

\[(5x+5)^2=100\]\[5x+5=\pm10\]

OpenStudy (anonymous):

OOh crap forgot about that! Whenever you square root don't forget to put plus or minus!

OpenStudy (anonymous):

my answers are 0,1? 1,4? 14 -14? 1, -4?

OpenStudy (anonymous):

5x + 5 = ±10 5x = -5 ± 10 x = -1 ± 2 x = 1 or x = -3 Not sure why it's not in your answers

OpenStudy (anonymous):

when 5x+5=+10, then x=1 when 5x+5=-10, then x=-3

OpenStudy (anonymous):

you guys are right... im soo sorry, i was looking at another problem.... thank u.. my bad.

OpenStudy (anonymous):

can you help me with, (2x+3) (exp 2) =25

OpenStudy (anonymous):

Same process, square root both sides first!

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

You get 2x+3 = ±5

OpenStudy (anonymous):

ok yeah